# Tangling-Untangling Cycle for Efficient Learning

### v1

@@ Abstract
The conventional wisdom of manifold learning is based on nonlinear dimensionality reduction techniques such as IsoMAP and locally linear embedding (LLE). We challenge this paradigm by exploiting the blessing of dimensionality. Our intuition is simple: it is easier to untangle a low-dimensional manifold in a higher-dimensional space due to its vastness, as guaranteed by Whitney embedding theorem. A new insight brought by this work is to introduce class labels as the context variables in the lifted higher-dimensional space (so supervised learning becomes unsupervised learning). We rigorously show that manifold untangling leads to linearly separable classifiers in the lifted space. To correct the inevitable overfitting, we consider the dual process of manifold untangling - tangling or aliasing- which is important for generalization. Using context as the bonding element, we construct a pair of manifold untangling and tangling operators, known as tangling-untangling cycle (TUC). Untangling operator maps context-independent representations (CIR) in low-dimensional space to context-dependent representations (CDR) in high-dimensional space by inducing context as hidden variables. The tangling operator maps CDR back to CIR by a simple integral transformation for invariance and generalization. We also present the hierarchical extensions of TUC based on the Cartesian product and the fractal geometry. Despite the conceptual simplicity, TUC admits a biologically plausible and energy-efficient implementation based on the time-locking behavior of polychronization neural groups (PNG) and sleep-wake cycle (SWC). The TUC-based theory applies to the computational modeling of various cognitive functions by hippocampal-neocortical systems.
I Introduction
A key challenge in machine learning (ML) is balancing specification and generalization (a.k.a. over-fitting and under-fitting). We want a model that generalizes well (but not so simplistic that it underfits) and is well-specified to capture the essential patterns in the training data (but not so complex that it overfits). This subtle balance is often achieved through statistical techniques like cross-validation, regularization, and choosing the right model complexity. From a geometric perspective, the challenge is intrinsically related to the manifold structure of the data acquired from the physical world. To discover and exploit such manifold structure, nonlinearity is commonly believed to be essential, as witnessed by kernel trick in support vector machine [48] and nonlinear dimensionality reduction (e.g., IsoMAP [45] and locally linear embedding [40]). However, nonlinearity is often at odds with rigorous proof and the simplicity criterion in mathematical and scientific research.
In this paper, we draw inspiration from the discovery of place cells [34] and its related aliasing problem [29, 50]. Nature has discovered an elegant solution to efficient learning through the evolution of mammalian brains, but its secret has remained elusive. We shed some light on this greatest puzzle by revisiting the role of specification and generalization from a manifold untangling/tangling perspective [27]. Whitney’s embedding theorem in topology [31] has shown that it is easier to untangle a manifold in a high dimensional space. Such intuition inspires us to introduce context dependency as the clue for manifold untangling. Using class labels C as contextual variables, we can show an arbitrary manifold $P(\textbf{X})$ can be untangled (admitting linear dichotomy [6]) in the lifted space $P(\textbf{X},\textbf{C})$. The opposite direction (tangling operator) is a simple integral transform that collapses/aliases all variables sharing the same context. By associating tangling/untangling with context-independent representation (CIR) and context-dependent representation (CDR), we rigorously show how to strike an optimal tradeoff between specificity and generalization by a pair of tangling and untangling operators, which we call tangling-untangling cycle (TUC).
Using TUC as the building block, we can construct more sophisticated learning algorithms based on Cartesian products and fractal geometry [28]. It can be shown that the Cartesian product of TUC is still a TUC, which supports the multiscale extension from low-dimensional to high-dimensional space. To mirror the nested structure of data acquired from the physical world, fractal-inspired extension is constructed where the inter-scale context dependency is induced by an index variable (a.k.a. positioning trick [26]). The efficiency of the proposed learning algorithms is supported by their biological implementations based on polychronization neural groups (PNG) [24]. The implementation of TUC by sleep-wake cycle [13] finishes our journey at its starting point - the discovery of place cells in the hippocampus [34]. The presence and absence of external stimuli (i.e., default state vs. perturbation state) offer an energy-efficient and biologically plausible implementation of untangling and tangling operators in wake and sleep modes, respectively. Overall, the main contributions of this paper can be summarized as follows:
  • A geometric insight about the blessing and curse of dimensionality leads to a new understanding of specification and generalization in ML, which nicely fits the emerging direct-fit-to-nature paradigm.
@@ I Introduction
  • A construction of manifold tangling-untangling cycle (TUC) for efficiency learning. Our construction represents a novel direct-fit-to-nature approach to unsupervised learning. Tangling and untangling operators are implemented by context dependency and integral transform, respectively.
@@
  • A biological connection of TUC with sleep-wake cycle (SWC) and its related cognitive functions. Untangling transforms CIR to CDR for specification in the wake mode; and untangling maps CDR back to CIR for generalization in the sleep mode.
@@
  • An application of TUC theory to model sensorimotor and social interactions. We use biological findings in various studies of cognitive functions as experimental support for the new theoretical framework.
II Geometric Intuition
Figure 1: Kernel trick achieves linear separability by mapping input data to a higher-dimensional feature space. $\phi:(x,y)\rightarrow(x,y,c)$, where $c=0,1$ denote red and blue classes, is a simple lifting idea that can be generalizable to a data manifold in arbitrary dimensions.
@@ II Geometric Intuition
In graph theory, there is a well-known educational puzzle named “Three Utilities Problem”. This puzzle can be formalized as a problem in topological graph theory by asking whether the complete bipartite graph $K_{3,3}$, has a graph embedding in the plane. The impossibility of the puzzle corresponds to the fact that $K_{3,3}$ is not a planar graph. However, such a puzzle can be easily solved if the planarity constraint is removed. This is a simple example illustrating the blessing of dimensionality to the problem of manifold untangling. In topology, Whitney Embedding Theorem [31] states that any smooth manifold can be embedded into a Euclidean space of double its dimension. Based on this classical result, we can prove the following lemma:
Lemma (Linear separability in high-dimensional space)
@@
Given two tangled manifolds $A,B\in R^{n}$, one can always linearly separate them in a high-dimensional space.
Proof. Using Whitney’s embedding theorem, we can also embed $A,B$ to a Euclidean space in $\tilde{A},\tilde{B}\in R^{2n}$. Then it is trivial to linearly separate $\tilde{A},\tilde{B}$ in $R^{4n}$.
@@
In machine learning (ML), the kernel trick used by support vector machine [42] represents a good example for illustrating the blessing of dimensionality. As shown in Fig. 1, linearly non-separable data sets can become untangled (linearly separable) after being mapped to a high dimensional feature space. The construction or learning of features has been at the foundation of all existing models, including both biological and artificial neural networks. Recent direct-fit models suggest that we can solve the manifold untangling problem without feature engineering [20]. This view is also consistent with intelligence without representation [3] and supported by the discovery of concept cells [38, 37]. As summarized in [9], identity-related feature information of an object manifold can be simultaneously conveyed by the gnostic units in the IT cortex along with contextual information (e.g., position, pose, and scale), implying a high-dimensional latent space.
III Context Dependency and Integral Transform for Untangling and Tangling
@@
Our intuition is based on observing the duality between specification (overfitting) and generalization (underfitting) in learning. An alternative approach to over/under-fitting is direct-fit-to-nature [20], which greatly inspired our approach in this work. Geometrically, we interpret overfitting/underfitting as manifold untangling and tangling, as shown in Fig. 2. Assuming sensory raw inputs are data belonging to a low-dimensional manifold embedded in a high-dimensional space, the objective of manifold untangling is to overfit the data for specification. This objective can be facilitated by embedding the manifold into a high-dimensional space, as Whitney’s embedding theorem suggests. Therefore, manifold untangling maps sensory raw input to a high-dimensional latent space where the fitting task becomes trivial (e.g., by linear dichotomy [6]). The dual process of manifold tangling is to underfit the data for generalization. A common example is aliasing [29], which reduces the dimensionality by collapsing the variables with a common state. We suggest an integral transform as a more general strategy than aliasing for generalization. In this section, we will study the pair of manifold untangling and tangling separately.
Figure 2: The tangling-untangling cycle (TUC) for solving overfitting/underfitting problems in efficient ML. Note its biological plausibility is consistent with sleep-wake cycle (SWC).
III-A Manifold Untangling: Dimensionality Expansion via Context Dependency
@@
The intuition behind context dependency can be easily explained by the toy example of binary classification (e.g., cats vs. dogs). Suppose we collect sensory raw data associated with two classes $\textbf{X}_{0},\textbf{X}_{1}\in R^{n}$ corresponding to classes 0 and 1, respectively. The raw data in $R^{n}$ is tangled (e.g., blue vs. red dots as shown in Fig. 1) but they can be easily untangled by lifting the dimensionality by one - i.e., $(\textbf{X}_{c},c)\in R^{n+1}$ where $c=0,1$ (a linear classifier $c=1/2$ is sufficient). Such an idea of dimensionality expansion can be better articulated as a strategy of context dependency where the label is explicitly coupled with the training data as a context variable. Note that the introduction of label-based context dependency transforms supervised learning to unsupervised learning [2] by direct-fit-to-nature [20]. We can generalize the strategy of context dependency for high-dimensional labels/context as follows.
Proposition 1 (Specification). Context Dependency for Manifold Untangling.
@@
Given sensory raw data $\textbf{X}\in R^{n}$ and its associated label $\textbf{C}\in R^{k}$, we can untangle the tangled data distribution $P(\textbf{X})$ by introducing context dependency as a joint PDF $P(\textbf{X},\textbf{C})\in R^{n+k}$.
@@
The proof of proposition 1 is a straightforward extension of the strategy used in binary classification. By considering the discrete representation of label or context C, linear separability is guaranteed in the expanded higher-dimensional space.
@@
III-B Manifold Tangling for Generalization: Dimensionality Reduction via Integral Transform
The dual operation of manifold tangling can be understood as the extension of aliasing [29]. The basic idea behind aliasing is to provide useful invariants by representing as equivalent states in which the same action/class is needed. Aliasing is good for generalization but has the risk of over-generalization. Such risk can be well managed by applying aliasing or its extended tangling operator to the untangled data in the high-dimensional space. The trick is to selectively remove or collapse hidden states where the context dependency dictates the criterion of selectivity. Using contextual information C as the guardrail, we are guaranteed a faithful generalization without the risk of over-generalization.
Proposition 2 (Generalization). Integral Transform for Manifold Tangling.
@@ III-B Manifold Tangling for Generalization: Dimensionality Reduction via Integral Transform
The pairing of manifold untangling and tangling operators forms a complete (TUC) cycle, as shown in Fig. 2. The connection of TUC with direct-fit models [20] is appealing - the objective of learning is not human-interpretable rules or world representations (i.e., “intelligence without representation” [3]). Instead, evolution has invented a localized computation for embodied cognition to interpolate over task-relevant manifolds (we will elaborate on this issue later when studying sensorimotor interaction). Manifold untangling offers a constructive solution to obtain over-parameterized models that are simple and parsimonious. More importantly, the connection of TUC with the evolutionary process is so powerful that we strongly believe it offers an important clue to unlock hidden secrets in developmental and ecological psychology. In summary, TUC forms the building block for our solution to efficient learning problems in general. Before we study the application of this new theory into neuroscience, we report several hierarchical extensions of TUC into chains and fundamental group of torus [21] to enrich our weaponry of constructing over-parameterized direct-fit models.
IV Hierarchical Extension and Connection with Fractal Geometry
First, we can spatially extend TUC with Cartesian products to form a fundamental group of the circles [21]. Such a dimensional increase is similar to the product PDF from the two marginals. Note that after manifold untangling, factorized linear dichotomies are conceptually similar to naive Bayes [39] (simple and parsimonious low-dimensional classifiers). Therefore, the Cartesian product preserves the independence of factorized distributions in the latent untangled space. In this manner, direct-fit models can be arbitrarily scaled to a high-dimensional space without suffering from the curse of dimensionality. Formally, we have the following result:

### v2

# On Computational Modeling of Sleep-Wake Cycle

@@ Abstract
Why do mammals need to sleep? Neuroscience treats sleep and wake as default and perturbation modes of the brain. It is hypothesized that the brain self-organizes neural activities without environmental inputs. This paper presents a new computational model of the sleep-wake cycle (SWC) for learning and memory. During the sleep mode, the memory consolidation by the thalamocortical system is abstracted by a disentangling operator that maps context-dependent representations (CDR) to context-independent representations (CIR) for generalization. Such a disentangling operator can be mathematically formalized by an integral transform that integrates the context variable from CDR. During the wake mode, the memory formation by the hippocampal-neocortical system is abstracted by an entangling operator from CIR to CDR where the context is introduced by physical motion. When designed as inductive bias, entangled CDR linearizes the problem of unsupervised learning for sensory memory by direct-fit. The concatenation of disentangling and entangling operators forms a disentangling-entangling cycle (DEC) as the building block for sensorimotor learning. We also discuss the relationship of DEC and SWC to the perception-action cycle (PAC) for internal model learning and perceptual control theory for the ecological origin of natural languages.
1 Introduction
We spend one-third of our life in sleep [9]; infants often sleep over half of the time every day. Why? As the brain’s default mode, sleep is believed to play an important role in the health and functioning of mammalian brains, encompassing several vital functions. Sleep deprivation often harms cognitive functions, including learning and memory. It is well known that memory consolidation, a process in which episodic memory becomes semantic memory, occurs during sleep [74]. Rapid eye movement (REM) sleep is believed to be important for consolidating procedural memories, those related to skills and tasks performed automatically without conscious awareness [69]. However, little is known about the mechanism of sleep, not to mention the computational modeling of the sleep-wake cycle (SWC) in the open literature.
In this paper, we fill this gap by drawing inspiration from the discovery of place cells [55] and its related aliasing problem [46, 75]. The aliasing problem, arising from the locality constraint of the receptive field, is solved by memory consolidation during sleep. Context-dependent unidirectional place cells in episodic memory are transformed into context-independent omnidirectional place cells in semantic memory [9]. Such a procedure can be intuitively interpreted as a manifold disentanglement or dimensionality reduction for generalization. The example of place cells for spatial navigation inspires us to generalize the role of memory consolidation into other cognitive tasks, such as perception and action. More specifically, we introduce context [^1] Context is defined as a set of conditions in which an input is uniquely coupled to an output [9]. dependency as design bias [25] or inductive bias [3] for entanglement (associated with specification or overfitting).
Figure 1: Illustration of sleep-wake cycle (SWC): entanglement/tangling occurs during the wake time (perturbation state) and disentanglement/untangling occurs during the sleep time (default mode).
Using motion as contextual variables, we can show an arbitrary manifold $P(\textbf{X})$ can be entangled during the wake mode into the lifted space $P(\textbf{X},\textbf{C})$ as a linearization strategy for classification, which is closely related to the recently proposed direct-fit strategy [30]. The opposite direction (disentangling operator) is abstracted by a simple integral transform that collapses/aliases all variables, sharing the same context variables for generalization. By associating entangling/disentangling with context-independent representation (CIR) and context-dependent representation (CDR), we rigorously show how the sleep-wake cycle (SWC) [21] might computationally implement a disentanglement-entanglement cycle (DEC) as shown in Fig. 1. In wake mode, the sensory cortex encodes the external environment $P(\textbf{Y})$ in CDR $P(\textbf{X},\textbf{C})$ in the latent space for specification during bottom-up processing; the projection from CDR $P(\textbf{X},\textbf{C})$ to CIR $P(\textbf{X})$ characterizes the recurrent top-down processing based on error feedback. When external environment Y is absent in the sleep mode, memory consolidation transforms CDR $P(\textbf{X},\textbf{C})$ into CIR $P(\textbf{X})$ by virtual “replay” for generalization.
Using DEC as the building block, we can construct more sophisticated sensorimotor learning algorithms for computational modeling of hippocampal-neocortical systems [8, 41]. In the framework of sensorimotor learning in the latent space [67], we present a computational model of the perception-action cycle (PAC) [22] based on the DECs of sensory and motor cortex, respectively. It is hypothesized that context dependency facilitates the task of sensorimotor learning in the latent space by linearizing error feedback in a task-optimal manner [54]. This hypothesis is supported by the duality between visual perception and motor control [57]. Similar to untangling in visual perception [16], disentanglement of actions into options has also been observed in motor control [77]. During sleep mode, PAC offers a proxy for modeling abstract thought (internal loop) in the absence of external stimuli; this internal loop can be extended to an external loop for social communication, which might be related to the ecological origin of natural languages [62].
Overall, the main contributions of this paper can be summarized as follows:
  • $\bullet$
    A geometric insight about the blessing and curse of dimensionality leads to a new understanding of inductive/design bias in ML - not all overfitting is bad, some are useful.
  • $\bullet$
@@ 1 Introduction
    A construction of disentanglement-entanglement cycle (DEC) for manifold learning. Our construction represents a novel direct-fit-to-nature approach to unsupervised learning. Tangling and untangling operators are implemented by context dependency and integral transform, respectively.
  • $\bullet$
@@
    A biological connection of DEC with sleep-wake cycle (SWC) and its related cognitive functions. Untangling transforms CIR to CDR for specification in the wake mode; and untangling maps CDR back to CIR for generalization in both wake (as top-down feedback) and sleep (as memory consolidation) mode.
  • $\bullet$
@@
    An application of DEC theory to model sensorimotor and hippocampal-neocortical interactions. We use biological findings in various studies of cognitive functions as experimental support for the new theoretical framework.
2 Geometric Intuition
@@
Graph theory has a well-known educational puzzle named “Three Utilities Problem". This puzzle can be formalized as a problem in topological graph theory by asking whether the complete bipartite graph $K_{3,3}$, has a graph embedding in the plane. The impossibility of the puzzle corresponds to the fact that $K_{3,3}$ is not a planar graph. However, removing the planarity constraint can easily solve such a puzzle. This is a simple example illustrating the blessing of dimensionality to the problem of manifold disentanglement/untangling. In our daily lives, the overpass at the intersection is another example of entanglement/tangling - the traffic jam problem can be effectively solved by lifting the space along the vertical dimension. By analogy, we can imagine that the pattern classification problem can be facilitated by a similar strategy of dimensionality increase (i.e., entanglement). We observe that such inductive bias can be exploited to discriminate patterns in a higher dimensional space by inducing motion as the contextual variable.
a)                                                                    b)
Figure 2: Comparison of different geometric intuitions: a) linear separability by kernel trick. b) linear separability by hyperplane separation theorem.
In topology, Whitney Embedding Theorem [48] states that any smooth manifold can be embedded into a Euclidean space of double its dimension. This fundamental result implies that an increase in dimensionality can facilitate the solution to pattern classification by linearizing the representation after embedding. We note that the geometric intuition of embedding differs from that of kernel trick (e.g., in support vector machine [73]), as shown in Fig. 2. Unlike kernel trick, embedding is independent of data representations; both can make nonlinear separability tractable by linear methods but only embedding can deliver data-agnostic solutions. In particular, we are interested in the class of linear dichotomies [12] as potential low-dimensional classifiers for learning task-state representations [54] after embedding. Based on Cover’s classical result [12], we can prove the following lemma related to the blessing of dimensionality:
Lemma 1 (Linear separability lemma)
@@
Given two entangled manifolds $A,B\in R^{n}$, one can always linearly separate them in a high-dimensional space.
Proof. In the hyperplane separation theorem [14], if both these sets are closed and at least one of them is compact, then there is a hyperplane in between them and even two parallel hyperplanes in between them separated by a gap (refer to Fig. 2). Using Whitney’s embedding theorem, we can also embed $A,B$ to a Euclidean space in $\tilde{A},\tilde{B}\in R^{2n}$. Then it is trivial to linearly separate $\tilde{A},\tilde{B}$ in $R^{4n+1}$ by constructing a hyperplane along the new dimension (e.g., consider the convex hulls of $\tilde{A},\tilde{B}$).
a)                                                         b)
Figure 3: Abstraction of SWC by DEC and its relationship to context dependency. a) The disentanglement-entanglement cycle (DEC) is used to solve overfitting/underfitting problems in manifold learning, whose biological plausibility is supported by the sleep-wake cycle (SWC). b) Implementation of DEC by a hysteresis loop (stable states “0” and “1” correspond to CIR and CDR, respectively). Note that trajectories “A” and “B” uniquely determine the context dependency (i.e., symmetry-breaking) between stimulus (S) and response (R).
@@
In machine learning (ML), the kernel trick used by support vector machine [65] represents a good example for illustrating the blessing of dimensionality. Based on Mercer’s theorem, linearly non-separable data sets can become untangled (linearly separable) after being mapped to a high-dimensional feature space. The construction or learning of features has been at the foundation of many existing learning models, including biological and artificial neural networks. Recent direct-fit-to-nature strategy [30] suggests that we can solve the manifold untangling problem without feature engineering, such as the construction of kernel machines [73]. The key idea is to assume a dense sampling of data manifold so generalization can be achieved by local interpolation instead of extrapolation on over-parameterized models. This view is also consistent with intelligence without representation [5] (i.e., “world as its own model”) and supported by the discovery of concept cells [59, 58]. As summarized in [16], identity-related feature information of an object manifold can be simultaneously conveyed by the gnostic units or neural population geometry [11] in the IT cortex along with contextual information (e.g., position, pose, and scale), implying a high-dimensional latent space in which all decisions can be made by a collection of low-dimensional linear dichotomies [12].
3 Entanglement and Disentanglement for Wake and Sleep
Our biological intuition is based on observing the alternating default and perturbation modes in SWC. It has been hypothesized in [9] that “sleep may serve to cement or consolidate memories by replaying parts and details of the waking experience”. The process of memory consolidation is believed to occur during REM sleep [74], which transforms episodic memory into semantic memory. Episodic memories are context-dependent representations (CDR) in which context is specified by the spatiotemporal position associated with the animal’s behavior. Semantic memories are context-independent representation (CIR) as the consequence of generalization. Geometrically, $CIR\rightarrow CDR$ corresponds to dimensionality expansion because context is induced as design bias [25]; $CDR\rightarrow CIR$ is implemented by dimensionality reduction because aliasing or collapse of contextual variables leads to generalization [46].
@@ 3 Entanglement and Disentanglement for Wake and Sleep
Computationally, we abstract the pair of dimensionality expansion and reduction operators into a cycle between CIR and CDR, as shown in Fig. 3a. Assuming sensory raw inputs are data belonging to a low-dimensional manifold embedded in a high-dimensional space, the objective of manifold entanglement or tangling is to overfit the data for specification by introducing context as design/inductive bias [3]. This objective can be facilitated by embedding the manifold into a high-dimensional space, as Whitney’s embedding theorem suggests. Therefore, manifold entanglement/tangling maps sensory raw input to a high-dimensional latent space where the task of separating different objects can be facilitated by linearization (e.g., by linear dichotomy [12]). The dual process of manifold disentanglement/untangling is to reduce the dimensionality of data representation for generalization. A common example is aliasing [46], which collapses the variables with a common state (context). This section will study the pair of manifold disentanglement and entanglement separately.
3.1 Wake Mode for Specification: Dimensionality Expansion via Context Dependency
@@
The intuition behind context dependency can be easily explained by the toy example of binary classification in supervised learning (e.g., cats vs. dogs). Suppose we collect sensory raw data associated with two classes $\textbf{X}_{0},\textbf{X}_{1}\in R^{n}$ corresponding to classes 0 and 1, respectively. The raw data in $R^{n}$ is generally not linearly separable but they can be made linearly separable by lifting the data distribution to a new dimension - i.e., $(\textbf{X}_{c},c)\in R^{n+1}$ where $c=0,1$ (a linear classifier $c=1/2$ is sufficient). Such an idea of dimensionality expansion can be better articulated as a strategy of context dependency that transforms a supervised learning problem into an unsupervised learning problem [2]. The key idea is that the class labels associated with different objects can be explicitly introduced as inductive bias or context variables.
Alternatively, we note that the spatio-temporal locality of the object often enforces the context information in the physical world. In other words, the local receptive field of a mammalian brain can only pay attention to a single object in space at any time. This strategy of overfitting with design bias implements recently proposed “direct-fit-to-nature” [30], which relies on over-parameterized models to provide mindless generalization. The shared vision between our context dependency and direct-fit is to densely sample the problem space so that local heuristics (e.g., nearest neighbor or averaging) will be sufficient for situating novel observations within the context of past observations. This vision is also consistent with intelligence without representation [5] which advocates for learning without explicitly representing underlying rules and principles in a human-interpretable way [30]. We can generalize the strategy of direct-fit by context-dependency for high-dimensional labels/context as follows.
Proposition 1 (Specification). Context Dependency for Direct-Fit.
@@
Given sensory raw data $\textbf{X}\in R^{n}$ and its associated label $\textbf{C}\in R^{k}$, we can directly fit the data distribution $P(\textbf{X})$ by introducing context dependency as a joint PDF $P(\textbf{X},\textbf{C})\in R^{n+k}$.
@@
The proof of Proposition 1 is a straightforward extension of the strategy used in binary classification. By considering the discrete representation of label or context C, linear separability is guaranteed in the expanded higher-dimensional space. Note that when the dimensionality of C is higher than that of X, we will observe high-dimensional encoding space as envisioned in direct-fit [30], which turns an extrapolation problem into an interpolation problem.
@@
3.2 Sleep Mode for Generalization: Dimensionality Reduction via Integral Transform
The dual operation of sleep mode can be understood as the extension of aliasing [46] by memory consolidation. The basic idea behind aliasing is to provide useful invariants or generalizations by representing equivalent states in which the same action/class is needed. Aliasing is usually achieved by dimensionality reduction, which collapses the distribution with shared context variables. More generally, we suggest that generalization can be achieved by mapping from CDR to CIR using integral transforms. The trick is to selectively remove or collapse hidden states C where the context dependency dictates the criterion of selectivity. Using contextual information C as the guardrail, we are guaranteed a faithful generalization without the risk of over-generalization. Formally, we have the following results.
Proposition 2 (Generalization). Integral Transform for Interpolation-based Generalization.
@@ 3.2 Sleep Mode for Generalization: Dimensionality Reduction via Integral Transform
Note that integral transforms are intrinsically interpolation or averaging in high-dimensional space. From the perspective of distributed representation, [33], we can interpret context C as a distributed set of representing X. In distributed representations, interpolation-based generalization happens when a neural network learns to map inputs to a high-dimensional space where similar inputs are closer together and dissimilar inputs are farther apart. Integral transform offers one step beyond interpolation-based generalization [33] because all noise and interference in task-irrelevant dimensions will be filtered out for improved robustness.
The pairing of specification and generalization operators forms a complete model for SWC, as shown in Fig. 3a. The connection of SWC with direct-fit models [30] is appealing - the objective of learning is not human-interpretable rules or world representations but rational behaviors for survival and prosperity in the physical world. Toward this objective, evolution has invented a localized computation for embodied cognition to interpolate over task-relevant manifolds (we will elaborate on this issue later when studying sensorimotor interaction). Direct-fit implemenyed by SWC offers a constructive solution to obtain over-parameterized models that are simple and parsimonious. More importantly, the connection of SWC with the evolutionary process is so powerful that we strongly believe it offers an important clue to unlock hidden secrets in developmental and ecological psychology. In summary, SWC forms the building block for our solution to efficient learning problems in general. Before we study the application of this new theory into neuroscience, we report hierarchical extensions of SWC into chains and fundamental group of torus [31] to enrich our weaponry of constructing over-parameterized direct-fit models.
4 Hierarchical Extension: from Cycles to Torus and Chain
First, we can spatially extend SWC with Cartesian products to form a fundamental group of the circles [31]. Such a dimensional increase is similar to the product PDF from the two marginals. Note that after manifold untangling, factorized linear dichotomies are conceptually similar to naive Bayes [61] (simple and parsimonious low-dimensional classifiers). Therefore, the Cartesian product preserves the independence of factorized distributions in the latent untangled space. In this manner, direct-fit models can be arbitrarily scaled to a high-dimensional space without suffering from the curse of dimensionality. Formally, we have the following result: